Find the differential of y=x³-2x²+5


1
Expert's answer
2021-01-21T11:24:15-0500

"To\;find\;:\;Differential\;of\;y=x^3-2x^2+5.\\
Pre-requisite:\;\;\;\\ 
1)\;\dfrac{d}{dx}(c)=0,\;where\;c\;is\;constant.\\"


"2)\;\dfrac{d}{dx}(cf(x))=\;c\dfrac{d}{dx}(f(x))"


"3)\;\dfrac{d}{dx}(x^n)=nx^{n-1},\;\;where\;n \in Z."


"\therefore \dfrac{d}{dx}(y)=\dfrac{d}{dx}(x^3-2x^2+5)"


"\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\dfrac{d}{dx}(x^3)-\dfrac{d}{dx}(2x^2)+\dfrac{d}{dx}(5)"


"\;\;\;\;\;\;\;\;\;\;\;\;\;\;=3(x^{3-1})-2(2)(x^{2-1})+0"


"\;\;\;\;\;\;\;\;\;\;\;\;\;\;=3x^{2}-4x"


"\therefore \dfrac{dy}{dx}=3x^2-4x"


"\therefore The \; differential\; has\;the\; following\; form :
dy=(3x^2-4x)dx"